# A Resolution Independent Neural Operator

### v2

@@ 2.1.1 Operator Learning with Predefined Trunk
where $\bar{\boldsymbol{\gamma}}^{(i)}_{\boldsymbol{\theta}}=\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}(\boldsymbol{\alpha}^{(i)})$, $\boldsymbol{\gamma}^{(i)}=\text{Proj}_{\Psi_{S}}(\mathcal{D}^{(i)}_{s})$, $\boldsymbol{\alpha}^{(i)}=\text{Proj}_{\Psi_{U}}(\mathcal{D}^{(i)}_{u})$, and $\tau$ is a penalty factor that controls the extent to which the second term is enforced. The first term in this objective function corresponds to the prediction part of the operator network, while the second term helps the branch network produce a representation consistent with the coefficients of the output basis functions. The reconstruction operator for a dictionary $\boldsymbol{\Psi}$ is defined as follows:
@@ 2.2 Dictionary Learning
In this section, we introduce two dictionary learning algorithms to identify a set of basis functions parameterized by neural networks, which are used to approximate signals defined on point cloud data. The first algorithm aims to adaptively find basis functions that capture the expected value of the residual across batches of signal realizations, while the second algorithm aims to find a basis each time that is needed to approximate only one realization at a time.
@@ 2.2.1 Batch-wise Learning
Figure 3: Dictionary learning (top) in the classical vector space setup and (bottom) in our proposed function space setup. Our method can be directly applied to point cloud signals that are sampled arbitrarily (but sufficiently richly) in terms of the number and location of discretized points. In our method, classical discrete atoms (basis vectors) are replaced by continuous, fully differentiable functions parameterized by neural networks.
Our goal here is to learn a dictionary that can represent the input function $u(x)$; however, the same procedure can be applied to the output function as well. In a general setup, we are given a dataset of $N$ recorded signals, denoted by $\mathcal{D}_{u}=\{\mathcal{D}_{u}^{(i)}\}_{i=1}^{N}$, where each element $\mathcal{D}_{u}^{(i)}$ contains an arbitrarily discretized version of the $i$-th function realization $u^{(i)}$, as described in the previous section and depicted in the bottom of Figure 3. If the number and locations of sensors (discretization points) are identical for all signals, the data setup resembles the classical PCA setup.
@@
where $\boldsymbol{\theta}$ concatenates the parameters of all neural network basis functions. Now, the challenge is to learn the parameterized dictionary $\boldsymbol{\Psi}(x;\boldsymbol{\theta})$ such that its atoms are orthogonal to each other.
@@
The projection operator $\text{Proj}_{\Psi}(\cdot):\mathcal{U}\mapsto\mathbb{R}^{|\Psi|}$ is defined with respect to the dictionary $\boldsymbol{\Psi}$ with $|\Psi|$ number of atoms, which returns the coordinates of the projection of the input function $u(x)$ onto the subspace spanned by the dictionary’s basis functions, such that an error measure (in the empirical risk minimization sense) between the original signal and its projection is minimized, as follows,
@@
where $\boldsymbol{\Psi}^{(i)}(\boldsymbol{x};\boldsymbol{\theta})$ represents the basis function values at the discretization points of the $i$-th realization $\{\boldsymbol{x}^{(i,j)}\}_{j=1}^{M^{(i)}}$, and $\boldsymbol{U}^{(i)}$ are the $i$-th realization values at those corresponding points $\{\boldsymbol{u}^{(i,j)}\}_{j=1}^{M^{(i)}}$. Notice that the kernel of the linear system scales with the dictionary size and not the discretization resolution. Therefore, this implementation is usually more computationally favorable than Eq. (13) which requires numerical integration and scales with the resolution size, particularly in cases where the number of basis functions is much less than the discretization resolution, which is the case in our numerical examples.
We have studied the ability of the proposed algorithm to find appropriate orthogonal basis functions in Appendix F, where we compare our algorithm with classical Gappy PCA or GPOD [13] in identifying basis functions from masked, incomplete data. Moreover, the possibility of using random basis functions is also explored and discussed.
@@ 2.2.2 Sample-wise Learning
Discussion: Normalization – From a theoretical perspective, the introduced algorithms aim to (weakly) enforce orthogonality of the identified basis functions. However, for several reasons, one might be interested in basis functions that are normalized as well, i.e., orthonormal basis functions. One reason could be attributed to potential identifiability issues associated with Eq. (11), where both the basis functions and coefficients are unknown. If one scales up the coefficients by a factor and simultaneously scales down the basis functions by the same factor, the same representation can be achieved, although the unknowns themselves have changed. The ADMM method mitigates this issue by solving for each unknown separately while fixing the others, thereby preventing simultaneous changes. However, controlling the scale or norm of either set of unknowns may further suppress such potential issues. Also, the relative scale of each basis function compared to others may impact the scale of its corresponding optimal coefficient relative to the coefficients of other basis functions. Since we may later use the basis coefficients in other downstream tasks (here, operator learning), it may be helpful to have learned bases with relatively the same scaling or norm. This way, the scale of their corresponding coefficients will have a particular meaning in terms of their relative importance in the prediction (coefficients with higher values indicate that their corresponding bases have more impact in the prediction). Several strategies are possible for the normalization of the learned basis functions: (1) one may use a normalization layer in the output layer of the neural network during the training to adjust the basis functions according to the batch statistics. In this study, we used an $L_{2}$-norm normalization layer of this type, where the normalized basis is $\psi_{\text{norm}}(\boldsymbol{x};\boldsymbol{\theta})=\psi(\boldsymbol{x};\boldsymbol{\theta})/\sqrt{\mathbb{E}_{x}[\psi^{2}(\boldsymbol{x};\boldsymbol{\theta})]}$. This type of normalization shares similarities with other batch-wise normalization layers commonly used in ML [1, 49]. (2) If the batch statistics of the given data do not accurately represent the actual statistics, one may normalize the learned basis after its training. To do this, a sufficiently large number of points can be artificially generated to empirically estimate the norm of the current un-normalized basis function, using a Markov Chain Monte Carlo (MCMC) estimate. This scaling factor can then be saved for later use before proceeding to learn the next basis function. In this work, we used the first approach in our numerical examples unless otherwise specified.
2.3 Parameterization via Implicit Neural Representation
@@
Use cases of INRs in the literature mostly deal with modeling the signal itself directly. However, in this work, instead of applying the INRs directly to the signal $u(\boldsymbol{x})$, we apply them to the basis functions $\psi_{l}(\boldsymbol{x})$, which serve as templates for modeling $u(\boldsymbol{x})$. One challenge in utilizing INRs is that conventional MLPs, despite their universal approximation property [18], struggle to learn fine details or high-frequency content due to a phenomenon known as spectral bias [30]. Several recent works aim to address this issue by modifying the conventional MLP architecture. Two prominent approaches include using sinusoidal activation functions, known as the SIREN approach [36], and incorporating random Fourier features in the first layer of the MLP, known as positional encoding [26, 41, 31]. Here, we use the SIREN approach to parameterize each newly added basis function $\psi_{l}(x;\boldsymbol{\theta}_{l})$ in Algorithm 1 due to its simplicity and the extensive success reported in various computer vision tasks [36, 8, 47].
@@
where $\boldsymbol{h}^{(l)}_{m}$ denotes the $m$-th hidden state, with $h_{0}\triangleq\boldsymbol{x}$ and $0\leq m\leq n$. $\boldsymbol{W}^{(l)}_{m}$ and $\boldsymbol{\beta}^{(l)}_{m}$ represent the typical weights and biases of the MLP, respectively. The vector $\boldsymbol{\theta}_{l}$ concatenates all the trainable parameters of the SIREN, i.e., $\boldsymbol{\theta}_{l}\equiv\left\{\boldsymbol{W}^{(l)}_{m},\boldsymbol{\beta}^{(l)}_{m}\right\}_{m=1}^{n+1}$. The convergence and accuracy of the SIREN depend on the appropriate initialization of the trainable parameters. Following the method proposed by [36], we initialize the weights drawn from a uniform distribution $\boldsymbol{W}_{m}\sim\text{Uniform}[-\omega_{0}\sqrt{6/\text{dim}(\boldsymbol{h}_{m-1})},\omega_{0}\sqrt{6/\text{dim}(\boldsymbol{h}_{m-1})}]$ where $\omega_{0}=1$ for all layers except the first layer, in which it is a hyperparameter that depends on the problem at hand. This ensures that the pre-activations are normally distributed with unit variance, leading to more stable training by preventing vanishing or exploding gradients. The biases are set to zero during the initialization stage.
@@
Although INRs could be more memory efficient per instance (e.g., an image) signal compression, their vanilla use case may not seem ideal when dealing with many instances. This is because each instance requires a separate INR with different parameters to be trained from scratch, without leveraging any information obtained from other instances or any potential common structures across instances. Notice that each training of an INR involves solving a non-convex optimization problem, which is typical for deep neural network training. One approach currently under exploration is to use meta-learning for better initialization of each INR, leading to fewer iterations, or some form of parameter sharing across instances [35, 40]. However, as explained, in our approach we do not directly model signals with INRs. Instead, we learn a dictionary of basis functions as INRs in an offline manner, which indirectly captures the common structure across instances. Then, at inference or for an unseen signal, we only solve a convex optimization problem to find the coefficients of the INR basis. This eliminates the need to introduce or train a new INR from scratch. In that regard, using INRs as basis functions may offer new ways to reduce the computational burden during inference when dealing with implicit representations of signals; however, this is not the focus of the current work.
@@ 3 Numerical Examples
In this section, we showcase the resolution independence of RINO by solving four numerical examples, where the data is synthesized from PDE solutions under different setups, ranging from 1D operators to operators with different supports in their input and output functions. The setup for the data generation and operator learning task is specified for each problem separately in their respective sections. Examples in Sections 3.1.1, 3.1.2, 3.1.3 and 3.1.4 are chosen from the literature or designed to verify the resolution independence of RINO with respect to the input function data, which is the main focus of this study. Example in Section 3.2 is designed to demonstrate the application of RINO in a more challenging scenario where the data realizations are obtained from finite element simulations of varying resolutions. The compactness and consistency of the vectorial representation achieved through our dictionary learning algorithm are discussed and studied in comparison with random projections in Appendix  F.
In cases where the PDE solver for data generation works on a fixed discretization, we randomly select only a subset of input function sensors for training to showcase RINO’s ability to handle irregular data. The locations and numbers of these sensors vary from sample to sample. Assuming an input function realization is discretized at $M$ points, a corresponding “random" realization with random discretization is identified by randomly selecting $M_{\text{rand}}$ points without replacement, where $M_{\text{rand}}$ is independently drawn for each realization from the uniform distribution $M_{\text{rand}}\sim\text{Uniform}[M_{\text{min}},M_{\text{max}}]$, and $0<M_{\text{min}}<M_{\text{max}}\leq M$.
@@ 3.1.1 Example 1: Antiderivative
where $u(x)$ is modeled as a Gaussian random process with zero mean and a covariance function defined by the radial basis function (RBF) with length scale $l=0.2$, i.e., $u(x)\sim\mathcal{GP}(0,\text{Cov}(x_{1},x_{2}))$; details are provided in Appendix G. We use the open-source dataset from [22], which includes 150 pairs of input-output functions for the training set and 1000 pairs for the test set. In the original dataset, the functions are discretized at 100 equally spaced sensor points. Following the procedure described earlier, we set $M_{\text{min}}=10$ and $M_{\text{max}}=60$ to generate random point clouds for the training and testing cases. In other words, each training data function has a randomly subsampled discretization of between 10 and 60 points.
@@
As described in the formulation section, the first step in RINO is to learn a dictionary to represent the input functions. This dictionary will be used to find embedding coordinates for the operator learning task, specifically for use in the branch network. Figure 4(a) shows the reconstruction errors in Algorithm 1 during the ADAM [19] iterations. Each red cross on the curve indicates the step at which a new basis function is added (line 8 in Algorithm 1). With a total of 9 learned basis functions, the data can be reconstructed with high accuracy, as shown by the distribution of errors in Figure 4(b) for both training and testing datasets. The learned basis functions are plotted in Figure 4(c).
@@
Figure 6: Antiderivative Example: RINO prediction for two queried input functions at inference and for the testing dataset: (a) Worst-case sample with the highest error in output function prediction, (b) Sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the queried input function data as ‘o’ and the masked points (not provided to the dictionary) as ‘x’, along with the reconstruction from dictionary learning. The second column shows the projection coefficients of the queried input function data onto the learned dictionary of basis function from Figure 4(c). The last column shows the output function prediction along with the true solution.
In Figure 6, the prediction of RINO at inference is shown for two queries from the test dataset. Figure 6(a) presents the results for the worst-case output function prediction. The first column shows the queried input function, which includes a point cloud with 35 random data points shown as ‘o’ along with the masked data points shown as ‘x’ that are not provided to the dictionary but are used to assess the accuracy of the reconstruction. The second column shows the projection coefficients of the input function data onto the learned dictionary, which has a total of 9 basis functions. The reconstruction using these coefficients is also plotted in the first figure. The last column shows the prediction of the output function along with the true model solution. Similarly, Figure 6(b) presents the same results for the test sample at the 25th percentile of the error distribution, meaning 75% of test cases have lower error than this sample. As the results show, the RINO method demonstrates good performance in providing predictions at inference for input function data recorded at an arbitrary (but sufficiently rich) number of sensors and locations.

### v3

@@ 2.1.1 Operator Learning Between Embedded Function Spaces (RINO)
where $\bar{\boldsymbol{\gamma}}^{(i)}_{\boldsymbol{\theta}}=\mathcal{F}_{\boldsymbol{\theta}}^{\text{ri}}(\boldsymbol{\alpha}^{(i)})$, $\boldsymbol{\gamma}^{(i)}=\text{Proj}_{\Psi_{S}}(\mathcal{D}^{(i)}_{s})$, $\boldsymbol{\alpha}^{(i)}=\text{Proj}_{\Psi_{U}}(\mathcal{D}^{(i)}_{u})$, and $\tau$ is a penalty factor that controls the extent to which the second term is enforced. Note that $\mathcal{F}^{\text{ri}}$ can be viewed as analogous to the branch net in DeepONet, however the architecture in Figure 2(b) is different in the sense that the neural operator in this setting is comprised of a single neural network operating between embedded function spaces. The first term in this objective function corresponds to the prediction part of the operator network, while the second term helps the branch network produce a representation consistent with the coefficients of the output basis functions. The reconstruction operator for a dictionary $\boldsymbol{\Psi}$ is defined as follows:
@@
A Discussion: If the second term of the training loss reaches zero, the first term will also achieve its minimum. This is because, prior to the operator learning step, the output basis functions are learned in a data-driven manner to ensure they can represent the output function within a specified tolerance. However, since the basis functions are not guaranteed to be perfectly orthogonal (as orthogonality is weakly enforced on finite sample points in a data-driven manner), there is a slight possibility of multiple sets of basis coefficients for a given realization producing the same minimal output reconstruction error. This underscores the importance of enforcing orthogonality, as it ensures the uniqueness of the embedding representation, thereby improving both stability and interpretability. In such cases, while the second term of the loss may have a higher error, the first term remains minimal. Consequently, the first term is the most critical, and the second term can be ignored. However, this simplification may lead to slower convergence during training. To address this (computational) challenge, one can pre-condition the operator network by initially enforcing the second term more strictly and gradually reducing its penalty weight as training progresses. Alternatively, one could bypass the first term entirely throughout the optimization process, depending on the specific training strategy, application requirements, and, importantly, the size of the embedding. This is because, in very high-dimensional spaces, the Euclidean distance between any two random vectors tends to become negligible, reducing the practical significance of the second term. However, in this work, since the learned basis functions are (weakly) orthogonal and the embedding dimensions do not exceed 100, we focus solely on the second loss term to expedite the operator learning training process.
@@ 2.2 Dictionary Learning
In this section, we introduce two new dictionary learning algorithms to identify a set of basis functions parameterized by neural networks, which are used to approximate signals defined on point cloud data. The first algorithm aims to adaptively find basis functions that capture the expected value of the residual across batches of signal realizations, while the second algorithm aims to find a basis each time one is needed to approximate only one realization at a time.
@@ 2.2.1 Batch-wise Learning
Figure 3: Dictionary learning (top) in the classical vector space setup and (bottom) in the proposed function space setup. The proposed method can be directly applied to point cloud signals that are sampled arbitrarily (but sufficiently richly) in terms of the number and location of discretized points. In the proposed method, classical discrete atoms (basis vectors) are replaced by continuous, fully differentiable functions parameterized by neural networks.
Throughout the following sections, we describe the procedure by learning a dictionary that can represent the input function $u(x)$; however, the same procedure is applied to the output function as well in the RINO. In a general setup, we are given a dataset of $N$ recorded signals, denoted by $\mathcal{D}_{u}=\{\mathcal{D}_{u}^{(i)}\}_{i=1}^{N}$, where each element $\mathcal{D}_{u}^{(i)}$ contains an arbitrarily discretized version of the $i$-th function realization $u^{(i)}$, as described in the previous section and depicted in the bottom of Figure 3. If the number and locations of sensors (discretization points) are identical for all signals, the data setup resembles the classical PCA setup.
@@
where $\boldsymbol{\theta}$ concatenates the parameters of all neural network basis functions. Now, the challenge is to learn the parameterized dictionary $\boldsymbol{\Psi}(x;\boldsymbol{\theta})$ such that its atoms are orthogonal to one another.
@@
The projection operator $\text{Proj}_{\Psi}(\cdot):\mathcal{U}\mapsto\mathbb{R}^{|\Psi|}$ is defined with respect to the dictionary $\boldsymbol{\Psi}$ with $|\Psi|=Q$ number of atoms, which returns the coordinates of the projection of the input function $u(x)$ onto the subspace spanned by the dictionary’s basis functions, such that an error measure (in the empirical risk minimization sense) between the original signal and its projection is minimized, as follows,
@@
where $\boldsymbol{\Psi}^{(i)}(\boldsymbol{x};\boldsymbol{\theta})$ represents the basis function values at the discretization points of the $i$-th realization $\{\boldsymbol{x}^{(i,j)}\}_{j=1}^{M^{(i)}}$, and $\boldsymbol{U}^{(i)}$ are the $i$-th realization values at those corresponding points $\{\boldsymbol{u}^{(i,j)}\}_{j=1}^{M^{(i)}}$. Notice that the kernel of the linear system scales with the dictionary size and not the discretization resolution. Therefore, this implementation is usually more computationally favorable than Eq. (13), which requires numerical integration and scales with the resolution size, particularly in cases where the number of basis functions is much less than the discretization resolution, which is the case in the numerical examples.
We have studied the ability of the proposed algorithm to find appropriate orthogonal basis functions in Appendix F, where we compare the proposed algorithm with classical Gappy PCA or GPOD [16] in identifying basis functions from masked, incomplete data. Moreover, the possibility of using random basis functions is also explored and discussed.
@@ 2.2.2 Sample-wise Learning
Discussion: Normalization – From a theoretical perspective, the introduced algorithms aim to (weakly) enforce orthogonality of the identified basis functions. However, for several reasons, one might be interested in basis functions that are normalized as well, i.e., orthonormal basis functions. One reason could be attributed to potential identifiability issues associated with Eq. (11), where both the basis functions and coefficients are unknown. If one scales up the coefficients by a factor and simultaneously scales down the basis functions by the same factor, the same representation can be achieved, although the unknowns themselves have changed. The ADMM method mitigates this issue by solving for each unknown separately while fixing the others, thereby preventing simultaneous changes. However, controlling the scale or norm of either set of unknowns may further suppress such potential issues. Also, the relative scale of each basis function compared to others may impact the scale of its corresponding optimal coefficient relative to the coefficients of other basis functions. Since we may later use the basis coefficients in other downstream tasks (here, operator learning), it may be helpful to have learned bases with relatively the same scaling or norm. This way, the scale of their corresponding coefficients will have a particular meaning in terms of their relative importance in the prediction (coefficients with higher values indicate that their corresponding bases have more impact in the prediction) which further enhances the interpretability of the proposed method. Several strategies are possible for the normalization of the learned basis functions: (1) one may use a normalization layer in the output layer of the neural network during the training to adjust the basis functions according to the batch statistics. In this study, we used an $L_{2}$-norm normalization layer of this type, where the normalized basis is $\psi_{\text{norm}}(\boldsymbol{x};\boldsymbol{\theta})=\psi(\boldsymbol{x};\boldsymbol{\theta})/\sqrt{\mathbb{E}_{x}[\psi^{2}(\boldsymbol{x};\boldsymbol{\theta})]}$. This type of normalization shares similarities with other batch-wise normalization layers commonly used in ML [1, 67]. (2) If the batch statistics of the given data do not accurately represent the actual statistics, one may normalize the learned basis after its training. To do this, a sufficiently large number of points can be artificially generated to empirically estimate the norm of the current un-normalized basis function, using a Monte Carlo estimate. This scaling factor can then be saved for later use before proceeding to learn the next basis function. In this work, we used the first approach in the numerical examples unless otherwise specified.
2.3 Parameterization via Implicit Neural Representation (Neural Field)
@@
Use cases of INRs or Neural fields in the literature mostly deal with modeling the signal itself directly. However, in this work, instead of applying the INRs directly to the signal $u(\boldsymbol{x})$, we apply them to the basis functions $\psi_{l}(\boldsymbol{x})$, which serve as templates for modeling $u(\boldsymbol{x})$. One challenge in utilizing INRs is that conventional MLPs, despite their universal approximation property [22], struggle to learn fine details or high-frequency content due to a phenomenon known as spectral bias [45]. Several recent works aim to address this issue by modifying the conventional MLP architecture. Two prominent approaches include using sinusoidal activation functions, known as the SIREN approach [53], and incorporating random Fourier features of different frequencies in the first layer of the MLP, known also as positional encoding for language modeling applications [38, 59, 46]. Here, we use the SIREN approach to parameterize each newly added basis function $\psi_{l}(x;\boldsymbol{\theta}_{l})$ in Algorithm 1 due to its simplicity and the extensive success reported in various computer vision tasks [53, 10, 65].
@@
where $\boldsymbol{h}^{(l)}_{m}$ denotes the $m$-th hidden state, with $h_{0}\triangleq\boldsymbol{x}$ and $0\leq m\leq n$. $\boldsymbol{W}^{(l)}_{m}$ and $\boldsymbol{\beta}^{(l)}_{m}$ represent the typical weights and biases of the MLP, respectively. The vector $\boldsymbol{\theta}_{l}$ concatenates all the trainable parameters of the SIREN, i.e., $\boldsymbol{\theta}_{l}\equiv\left\{\boldsymbol{W}^{(l)}_{m},\boldsymbol{\beta}^{(l)}_{m}\right\}_{m=1}^{n+1}$. The convergence and accuracy of the SIREN depend on the appropriate initialization of the trainable parameters. Following the method proposed by [53], we initialize the weights drawn from a uniform distribution $\boldsymbol{W}_{m}\sim\text{Uniform}[-\omega_{0}\sqrt{6/\text{dim}(\boldsymbol{h}_{m-1})},\omega_{0}\sqrt{6/\text{dim}(\boldsymbol{h}_{m-1})}]$ where the frequency parameter $\omega_{0}=1$ for all layers except the first layer, in which it is a hyperparameter that depends on the problem at hand. This ensures that the pre-activations are normally distributed with unit variance, leading to more stable training by preventing vanishing or exploding gradients. The biases are set to zero during the initialization stage.
@@
Although INRs could be more memory efficient per instance signal compression (e.g., an image), their vanilla use case may not seem ideal when dealing with many instances. This is because each instance requires a separate INR with different parameters to be trained from scratch, without leveraging any information obtained from other instances or any potential common structures across instances. Notice that each training of an INR involves solving a non-convex optimization problem, which is typical for deep neural network training. One approach currently under exploration is to use meta-learning for better initialization of each INR, leading to fewer iterations, or some form of parameter sharing across instances [52, 58, 15]. However, as explained, in the proposed approach we do not directly model signals with INRs. Instead, we learn a dictionary of basis functions as INRs in an offline manner, which indirectly captures the common structure across instances. Then, at inference or for an unseen signal, we only solve a convex optimization problem to find the coefficients of the INR basis. This eliminates the need to introduce or train a new INR from scratch. In that regard, using INRs as basis functions may offer new ways to reduce the computational burden during inference when dealing with implicit representations of signals; however, this is not the focus of the current work.
@@ 3 Numerical Examples
In this section, we showcase the resolution independence of RINO by solving several numerical examples, where the data is synthesized from PDE solutions under different setups, ranging from 1D operators to operators with different supports in their input and output functions. The setup for the data generation and operator learning task is specified for each problem separately in their respective sections. Examples in Sections 3.1.1, 3.1.2, 3.1.3 and 3.1.4 are chosen from the literature or designed to verify the resolution independence of the RI-DeepONet with respect to the input function data, since these simple functions do not require an advanced embedding of the output. Examples in Section 3.2 are designed to demonstrate the application of RINO in more challenging scenarios where the data realizations are obtained from finite element simulations of varying resolutions. In these cases, an embedding of the output is also beneficial. The compactness and consistency of the vectorial representation achieved through the proposed dictionary learning algorithm are discussed and studied in comparison with random projections in Appendix  F. Moreover, the generalization, interpolation, and extrapolation capabilities of the proposed dictionary learning approach are discussed there. Additional potential advantages of RINO over RI-DeepONet in terms of generalization are examined in Appendix G.
In cases where the PDE solver for data generation works on a fixed discretization, we randomly select only a subset of input function sensors for training to showcase the proposed method’s ability to handle irregular data. The locations and numbers of these sensors vary from sample to sample. Assuming an input function realization is discretized at $M$ points, a corresponding “random” realization with random discretization is identified by randomly selecting $M_{\text{rand}}$ points without replacement, where $M_{\text{rand}}$ is independently drawn for each realization from the uniform distribution $M_{\text{rand}}\sim\text{Uniform}[M_{\text{min}},M_{\text{max}}]$, and $0<M_{\text{min}}<M_{\text{max}}\leq M$.
@@ 3.1.1 Example 1: Antiderivative
where $u(x)$ is modeled as a Gaussian random process with zero mean and a covariance function defined by the radial basis function (RBF) with length scale $l=0.2$, i.e., $u(x)\sim\mathcal{GP}(0,\text{Cov}(x_{1},x_{2}))$; details are provided in Appendix H. We use the open-source dataset from [34], which includes 150 pairs of input-output functions for the training set and 1000 pairs for the test set. In the original dataset, the functions are discretized at 100 equally spaced sensor points. Following the procedure described earlier, we set $M_{\text{min}}=10$ and $M_{\text{max}}=60$ to generate random point clouds for the training and testing cases. In other words, each training data function has a randomly subsampled discretization of between 10 and 60 points.
@@
As described in the formulation section, the first step in the proposed approach is to learn a dictionary to represent the input functions (again, here we only encode the input). This dictionary will be used to find embedding coordinates for the operator learning task, specifically for use in the branch network of the RI-DeepONet. Figure 4(a) shows the reconstruction errors in Algorithm 1 during the ADAM [28] iterations. Each red cross on the curve indicates the step at which a new basis function is added (line 8 in Algorithm 1). With a total of 9 learned basis functions, the data can be reconstructed with high accuracy, as shown by the distribution of errors in Figure 4(b) for both training and testing datasets. The learned basis functions are plotted in Figure 4(c).
@@
Figure 6: Antiderivative Example: RI-DeepONet prediction for two queried input functions at inference and for the testing dataset: (a) Worst-case sample with the highest error in output function prediction, (b) Sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the queried input function data as ‘o’ along with the reconstruction from dictionary learning. The second column shows the projection coefficients of the queried input function data onto the learned dictionary of basis function from Figure 4(c). The last column shows the output function prediction along with the true solution.
In Figure 6, the prediction of RI-DeepONet at inference is shown for two queries from the test dataset. Figure 6(a) presents the results for the worst-case output function prediction. The first column shows the queried input function, which includes a point cloud with 35 random data points shown as ‘o’ along with the reconstructed input from dictionary learning and the true function. The second column shows the projection coefficients of the input function data onto the learned dictionary, which has a total of 9 basis functions. The last column shows the prediction of the output function along with the true model solution. Similarly, Figure 6(b) presents the same results for the test sample at the 25th percentile of the error distribution, meaning 75% of test cases have lower error than this sample. As the results show, the RI-DeepONet method demonstrates good performance in providing predictions at inference for input function data recorded at an arbitrary (but sufficiently rich) number of sensors and locations.